Your client has been given a trust fund valued at $1.57 million. She cannot access the money until she turns 65 years old, which is in 20 years. At that time, she can withdraw $23,500 per month. If the trust fund is invested at a 4.5 percent rate, how many months will it last your client once she starts to withdraw the money?
Answer #1

PV =1570000
Number of Years =20
PMT =23500
Rate =4.5%
FV at time of retirement =PV*(1+r)^n =1570000*(1+4.5%)^20 =3,786,391.0190

Monthly Withdrawal =23500
Rate per month =4.5%/12
Number of months using excel formula =NPER(4.5%/12,-23500,3786391.0190) =247.63 months